OpenStudy (anonymous):

Las Vegas Nevada, is 445 miles from Reno, Nevada. If the distance on the map is 11 1/8 inches, find the scale used to the map.

OpenStudy (anonymous):

convert 11 1/8 to miles and do 445/(that number)

OpenStudy (anonymous):

How do I convert that into miles. You mean convert it into an improper fraction?

OpenStudy (anonymous):

= 0.000175584

OpenStudy (anonymous):

Im not used to the US system.. I can only do calculus in the metric one.. I told you how to do it in my way. Im not sure if yours is the same. Maybe you should wait for someone thats used to inches

OpenStudy (anonymous):

Okay Thanks anyways :)

OpenStudy (anonymous):

lets give it another try 11.125 (i) = 445 (m) 445/11.125 = 40 1 inch = 40 miles

OpenStudy (anonymous):

I did this: 11.125 (i) = 445 (m) 445/11.125 = 40 1 inch = 40 miles

OpenStudy (anonymous):

Like yours!!

OpenStudy (anonymous):

GOod thank you! :)

OpenStudy (anonymous):

you're welcome :)